A standing wave on a string of length L = 3 m is described by the following equation: y(x,t) = 0.08 sin⁡(2πx) cos(300πt). The fundamental frequency, f1, is:


25


1
Expert's answer
2021-05-20T11:25:00-0400

Gives

Length (L)=3 m

"y(x,t)=0.08sin(2\pi x)cos(300\pi t)\rightarrow(1)"

Fundamental frequency "f_1=?"

"y(x,t)=Asin(k x)cos(wt)\rightarrow(2)"

We know that

"2L=n\lambda"

Put value

"2\times3=n\times1"

"n=6"

"K=\frac{2\pi}{\lambda}"

eqution (1)and(2)comperision

"2\pi=\frac{2\pi}{\lambda}"

"\lambda=1m"

Number of node =6

Wave length=1m

Fundamental frequency "f_1=\frac{\frac{w}{2\pi}}{n}"

Put value

"f_1=\frac{\frac{300\pi}{2\pi}}{6}"

"f_1=25"